A toy rocket is launched from the ground with an initial velocity of 80 feet per second. The height h of the rocket after t seconds is given by the equation h =80t-16t^2. How long after it is launched will the rocket be at it's highest point?
Answer by stanbon(48568)

A toy rocket is launched from the ground with an initial velocity of 80 feet per second. The height h of the rocket after t seconds is given by the equation h =80t-16t^2. How long after it is launched will the rocket be at it's highest point?
----------------
You have a quadratic with a = -16 and b=80
The maximum height occurs where t = -b/2a = -80/(2*-16) = 2.5 seconds
================
Cheers,
Stan H.